Louisiana State University Health Sciences Center Administration & Finance
 

Thursday, April 05, 2007

  Dental School :

  • LSUHSC Computer Services did some spot checks on the power supply for the upgrades that are planned for the network system in both the Clinic and Administration Buildings . 
    • They are also looking for a room to store 2700 ft of cable trays that are required for the upgrade/wiring project. 
  • Pflueger Electric reset the breaker on the switch gear for Chiller # 2 on the second floor in the Power Plant.  Each time they turned on the switch at the Chiller, it threw the main breaker.
    • Pflueger will contact the engineering firm to find out how to correct the problem.
  • Zimmer has completed the removal of ceiling grid on the 6th floor and will proceed to the 7th floor tomorrow.

 1542 Tulane:

  • Ordes has been working 10 hr. days Monday thru Thursday and 8 hrs. on Friday.  They are still running the 4 inch conduit. 
    • They are also preparing to install another electrical panel in the basement.
  • Stratos Elevator is working 8 hr. days.  They are still working on the interior of elevator #3 cab. 
    • They have removed and disposed of all of their old parts that were in the hall.
  • Arc Mechanical is working 10 hr. days Monday thru Thursday.  They are working on the steam station in the basement. 
    • Some of the replacement valves are not the same size as the old ones and some additional piping work is required.

 Irrigation Project: 

  • LSUHSC Facility Services personnel were able to put in 4 of the sleeves under sidewalks.  A quote on installing the driveway sleeves will be in next week.